# On the long-time limit of the mean curvature flow in closed manifolds

### v1

@@ Abstract.
In this article we show that generally almost regular flows, introduced by Bamler and Kleiner, in closed $3$–manifolds will either go extinct in finite time or flow to a collection of smooth embedded minimal surfaces, possibly with multiplicity. Using a perturbative argument then we construct piecewise almost regular flows which either go extinct in finite time or flow to a stable minimal surface, possibly with multiplicity. We apply these results to construct minimal surfaces in $3$–manifolds in a variety of circumstances, mainly novel from the point of the view that the arguments are via parabolic methods.
@@ 1. Introduction
The mean curvature flow, as the gradient flow of the area functional, is a natural tool to find minimal surfaces – in particular area minimizing and stable ones. The $1$-dimensional mean curvature flow in surfaces, known as the curve shortening flow, has been proven to be a very powerful tool for constructing geodesics. An especially important accomplishment was by Grayson [21], who used the curve shortening flow to settle Lusternik and Schnirelman’s proposal to construct $3$ distinct closed embedded geodesics in $(S^{2},g)$.
A central obstruction to using $n$–dimensional mean curvature flow in $(n+1)$–dimensional ambient manifold to construct minimal surfaces is the more complicated nature of singularity formation in dimensions $n\geq 2$, and weak notions of the flow are necessary. There are a number of reasonable definitions of weak flows to handle this issue; these are discussed in section 2. Very recently, Bamler and Kleiner [4] introduced the notion of almost regular flow, which are well-behaved Brakke flows whose so–called local scale functions satisfy a certain integrability condition. They proved that in ${\mathbb{R}}^{3}$ all generic flows (i.e. flows which only develop mean convex singularities) and hence, by an approximation argument, inner and outermost flows emanating from an embedded closed surface are almost regular. In this framework, they then proved the multiplicity one conjecture of Ilmanen [30]. In this article we use the notions of generic and almost regular flows in general $3$–manifolds and first show the following:
@@
Let $M$ be a $2$-sided properly embedded surface in a closed Riemannian $3$–manifold $(N,g)$, and let $M_{t}$ be an almost regular flow from $M$. Then either $M_{t}$ goes extinct in finite time or there exists a sequence $t_{i}\to\infty$ for which $M_{t_{i}}$ converges to $m_{1}\Sigma_{1}+\cdots+m_{k}\Sigma_{k}$ in the sense of varifolds, where each $\Sigma_{j}$ is a smoothly embedded minimal surface and each $m_{j}$ is a positive integer. Moreover, all the $\Sigma_{j}$’s are disjoint.
We point out that some of the $\Sigma_{j}$ may be nonorientable, even if the initial data is $2$–sided: consider for instance a situation where the boundary of the tubular neighborhood of an embedded projective plane $P$ (say in the 3 manifold $N={\mathbb{R}}P^{3}$), homeomorphic to $S^{2}$, flows back to $P$ as $t\to\infty$. This can also be interpreted as saying that the topology of the flow may actually "increase" in the limit; we discuss more precisely to what extent the topology of the $\Sigma_{i}$ can be controlled in terms of the initial data $M$ in Proposition 3.7 below. We note that apriori the limit in Theorem 1.1 may not be unique; if the ambient manifold is analytic and the convergence is with multiplicity one we may appeal to the Łojasiewicz-Simon inequality to see this is the case, but we assume/show neither in the statement above. However, we are able to show the limit stable minimal surface is unique whenever it is strictly stable. Recall below that a stable minimal surface is strictly stable if the first eigenvalue of the linearized operator is positive, and as pointed out for example in [48] this notion continues to make sense even for $1$-sided minimal surfaces:
@@
See Theorem 3.6 below. We would like to point out that the uniqueness of limit in geometric measure theory, in general, is quite hard when the multiplicity of convergence is greater than one.
Next, we further study what the possible limit minimal surfaces can be. Because the flow is the gradient of the area functional one expects to typically find area minimizing or at least stable minimal surfaces using it but nevertheless, in general, the long-time limit of mean curvature flow can be unstable. For curve shortening flow, a well-known example is the flow $\gamma_{t}$ of a closed embedded curve $\gamma\subset(S^{2},g_{round})$ which bisects the sphere. It will converge, as $t\to\infty$, smoothly to an equator of the sphere; curves $\gamma$ which do not bisect the sphere instead flow to round points. Most curves $\gamma$ do not bisect the sphere and we can always perturb one that does to one which does not, of course, so most curve shortening flows cannot converge to an unstable geodesic. This observation motivates us to consider to which extent we can show the flow will generically avoid unstable minimal surfaces.
The main tool we use to pursue this is the avoidance principle. In ${\mathbb{R}}^{n+1}$, if two mean curvature flows of closed hypersurfaces are disjoint at time 00, then they will remain disjoint for all the future time. Moreover, the distance between these two flows is increasing. This basic property motivates the definition of weak set flows by Ilmanen [28] and White [50]. The avoidance principle in a manifold is more subtle, because in general two disjoint mean curvature flows can actually get closer – consider for instance the flow of small graphs over strictly area minimizing minimal surfaces. In a manifold equipped with a positive Ricci curvature metric, the avoidance principle is very strong though and lets us easily show the following:
@@
In a Ricci positive manifold $(N^{n},g)$, the mean curvature flow of a generic embedded hypersurface $M$ will go extinct in finite time.
This can be viewed as a generalization of the generic behavior of the curve shortening flows in $S^{2}$. As a corollary this yields another proof of the classical fact that $H_{2}(N,{\mathbb{R}})$ for $3$–manifolds $N$ of positive Ricci curvature, see Corollary 4.1. In the case $(N,g)$ is not Ricci positive, there could be locally minimizing surfaces and we cannot expect the flows starting from generic embedded hypersurfaces to go extinct in finite time. However, we can use the avoidance principle to show that after perturbations, the flow can avoid those unstable limits. Our next theorem, inspired by work of Colding and Minicozzi [12], says that in closed $3$–manifolds one may construct, from any embedded initial data, piecewise generic mean curvature flows which avoid unstable minimal surfaces as limits:
@@
Let $M^{2}$ be a $2$–sided properly embedded surface of a closed compact Riemannian $3$–manifold $(N^{3},g)$. Then there exists a piecewise almost regular flow emanating from it which either goes extinct in finite time, or there exist $t_{i}\to\infty$, such that $M_{t_{i}}$ converges possibly with multiplicity to a (potentially disconnected) stable minimal surface.
When the ambient metric is bumpy, which is a generic condition by work of White [49, 54], we observe that by Theorem 1.2 the limit in the statement above is unique. Note that in light of [14] there are examples where the flow asymptotically converges to a stable limiting surface with multiplicity greater than one so higher multiplicity convergence in the above may actually occur, and in the case the limit is strictly stable seems to be robust under further perturbations. We will give the definition of piecewise almost regular flow in the section below but roughly speaking it is a concatenation of almost regular flows where by–hand isotopies are performed at the jumps; the number of perturbations by isotopy taken will be finite, and the perturbations can be taken as small as we wish. In that sense the statement above is a statement on generic flows following the convention of Colding and Minicozzi [12].
Of course, it is certainly desirable to only require perturbing the initial data. The obstruction to doing so in the statement above essentially boils down to the possibility of convergence to an unstable minimal surface with multiplicity; in analogy to singularity analysis and the resolution of the multiplicity one conjecture by Bamler and Kleiner [4] one might expect this to not occur; see the analogous situation in Almgren-Pitts minmax theory [57]. We recall the following conjecture in [14]:
Conjecture 1.5 (Conjecture 1.3 in [14]).
@@
Despite this, we point out that as applications we get flow proofs of the following existence theorems for stable minimal surfaces. In the proofs of these we will often specifically use piecewise generic flows, for instance because the topological change through singularities is easy to understand. We also recall that obtaining stability of the minimal surfaces is helpful because these often have better controlled topology and geometry, for instance in the context of positive scalar curvature. These corollaries are not new and there have been long standing proofs of them using geometric measure theory or harmonic map theory, see [19, 47, 43, 39, 38, 20], among others. On the other hand, the construction of these minimal surfaces via flows, along with a number of fairly explicit perturbations, is quite amenable to numerical approximation with the caveat that one has found appropriate initial data to start with. In the following statements $(N,g)$ is a compact orientable Riemannian $3$–manifold, where in this case oriented surfaces are $2$–sided and vice versa:
@@
Note that in the statement above we invoke the sphere theorem to find embedded initial data for which to start the flow and so the above does not constitute a new proof of it. As we pointed out already compared to other techniques a pitfall of using the mean curvature flow to find minimal surfaces is that it is best to start with well-prepared (i.e. embedded) initial data. In some cases the flow can produce stable minimal surfaces of nontrivial topology, the most basic of which is the following.
@@
Suppose $N$ has trivial second homotopy group and contains a homologically nontrivial embedded torus. Then $N$ contains an embedded stable minimal surface homeomorphic either to a torus or a Klein bottle.
Recall that hyperbolic manifolds are aspherical, so the space of such $3$–manifolds with $\pi_{2}$ trivial is quite large and is even generic amongst $3$–manifolds from a number of perspectives, for example see [16, 35]. For surfaces of more general topology, we can show the following result which is very much in the spirit of the existence result of Freedman, Hass, and Scott [20] – in fact it implies the result above as we discuss in section 5. The surfaces below are known as algebraically incompressible surfaces, and are incompressible in the more geometric sense via Dehn’s lemma:
@@
By double cover of an embedded surface $P\subset N$ here we mean specifically a surface which is the boundary of a tubular neighborhood of $P$ in $N$. It is a fact (see for instance Lemma 3.6 in [22]) that every class of $H_{2}(N,{\mathbb{Z}})$ can be represented by an orientable properly embedded surface $M$. Recalling the Thurston norm $||\sigma||_{T}$ of $\sigma\in H_{2}(N,{\mathbb{Z}})$ is the smallest attained genus of embedded surfaces homologous to $\sigma$, we see any such surface $M$ realizing the norm must be algebraically incompressible. If $||\sigma||_{T}\geq 1$ we can apply the Corollary above then, and clearly this must be the case if we suppose that $\pi_{2}(N)$ is trivial. We can record the discussion above as follows:
@@
Suppose $\sigma\in H_{2}(N,{\mathbb{Z}})$, with $\pi_{2}(N)$ trivial. Then either there exists an embedded oriented stable minimal surface $\Sigma$ in the class of $\sigma$ that realizes the Thurston norm of $\sigma$ or there is an element realizing the Thurston norm which is the double cover of a embedded nonoriented stable minimal surface.
Finally, we note that Theorem 1.1 lets us substitute the tightening process in minmax for mean curvature flow, at least in some cases. A major issue in applying the flow to a sweepout is that the flow of a sweepout may not remain a sweepout, but this can be compensated for if the flow of the elements of it will be nonfattening. In particular we show the following:
@@
Originally this result was proved by Simon and Smith using minmax theory, see also [9]. As to the actual best result known we point out that recently Haslhofer and Ketover in [25] showed there are at least two minimal spheres in $S^{3}$ with a bumpy metric, and very recently Zhichao Wang and Xin Zhou in [56] proved that there are at least $4$ embedded minimal spheres in $S^{3}$ equipped with either a bumpy or Ricci positive metric. We expect that if Conjecture 1.5 is true then the mean curvature flow can also be used to recover their results. We emphasize that in the above we are not applying Theorem 1.4 and the minimal surface above is potentially unstable, as must be the case when for instance $N$ is Ricci positive.
We conclude the introduction with a remark on the compactness assumptions made throughout. When the ambient manifold is not closed, but merely complete with uniform bounded curvature and injectivity radius lower bound, Theorems 1.1 and 1.4 remain valid, if we include the possibility that the flow can escape from any compact set. For example if $N$ is $S^{2}\times{\mathbb{R}}$ with an infinite long trumpet shape (i.e. a cusp) in one direction the flow of an appropriate $S^{2}$ section of it will escape to spatial infinity/clear out, neither terminating or converging to a minimal surface. It seems to be a more complicated manner yet to generalize our results to the case $M$ is allowed to be noncompact (say properly embedded in a noncompact $N$), one very basic reason being that the area can be infinite. As before the flow may clear out in the manner of the example above but it seems that in some cases if the flow doesn’t clear out or go extinct one could show it must flow to a minimal surface. For a basic example, a well known consequence of Ecker and Huisken [17] says that uniformly Lipschitz graphs of bounded height in ${\mathbb{R}}^{n}$ must converge to flat, hence minimal, planes.

### v2

@@ Abstract.
In this article, we show that generally almost regular flows, introduced by Bamler and Kleiner, in closed 33–manifolds will either go extinct in finite time or flow to a collection of smooth embedded minimal surfaces, possibly with multiplicity. Using a perturbative argument then we construct piecewise almost regular flows which either go extinct in finite time or flow to a stable minimal surface, possibly with multiplicity. We apply these results to construct minimal surfaces in 33–manifolds in a variety of circumstances, mainly novel from the point of the view that the arguments are via parabolic methods.
2000 Mathematics Subject Classification:
53E10, 53A10
@@ 1. Introduction
The mean curvature flow, as the gradient flow of the area functional, is a natural tool to find minimal surfaces – in particular area minimizing and stable ones. The $1$-dimensional mean curvature flow in surfaces, known as the curve shortening flow, has been proven to be a very powerful tool for constructing geodesics. An especially important accomplishment was by Grayson [Grayson89_CSF], who used the curve shortening flow to settle Lusternik and Schnirelman’s proposal to construct $3$ distinct closed embedded geodesics in $(S^{2},g)$.
A central obstruction to using nn–dimensional mean curvature flow in (n+1)(n+1)–dimensional ambient manifold to construct minimal surfaces is the more complicated nature of singularity formation in dimensions $n\geq 2$. Therefore, weak notions of the flow are necessary. There are a number of reasonable definitions of weak flows to handle this issue; these are discussed in section 2. Very recently, Bamler and Kleiner [BamlerKleiner23_Multiplicity1] introduced the notion of almost regular flow, which are well-behaved Brakke flows whose so–called local scale functions satisfy a certain integrability condition. They proved that in ${\mathbb{R}}^{3}$ all generic flows (i.e., flows which only develop mean convex singularities) and hence, by an approximation argument, inner and outermost flows emanating from an embedded closed surface are almost regular. In this framework, they then proved the Multiplicity One Conjecture of Ilmanen [Ilmanen95_Sing2D]. In this article, we use the notions of generic and almost regular flows in general 33–manifolds and first show the following:
@@
Let $M$ be a $2$-sided properly embedded surface in a closed Riemannian 33–manifold $(N,g)$, and let $M_{t}$ be an almost regular flow from $M$. Then either $M_{t}$ goes extinct in finite time or there exists a sequence $t_{i}\to\infty$ for which $M_{t_{i}}$ converges to $m_{1}\Sigma_{1}+\cdots+m_{k}\Sigma_{k}$ in the sense of varifolds, where each $\Sigma_{j}$ is a smoothly embedded minimal surface and each $m_{j}$ is a positive integer. Moreover, all the $\Sigma_{j}$’s are disjoint.
We point out that some of the $\Sigma_{j}$ may be nonorientable, even if the initial data is 22–sided: consider for instance a situation where the boundary of the tubular neighborhood of an embedded projective plane $P$ (say in the 3 manifold $N={\mathbb{R}}P^{3}$), homeomorphic to $S^{2}$, flows back to $P$ as $t\to\infty$. This can also be interpreted as saying that the topology of the flow may actually "increase" in the limit; we discuss more precisely to what extent the topology of the $\Sigma_{i}$ can be controlled in terms of the initial data $M$ in Proposition 3.7 below. We note that a priori the limit in Theorem 1.1 may not be unique; if the ambient manifold is analytic and the convergence is with multiplicity one, we may appeal to the Łojasiewicz-Simon inequality to see this is the case by [Simon83Asym], but we assume/show neither in the statement above. However, we are able to show that the limit stable minimal surface is unique whenever it is strictly stable. Recall below that a stable minimal surface is strictly stable if the first eigenvalue of the linearized operator is positive, and as pointed out, for example, in [Urbano13_one-sided], this notion continues to make sense even for $1$-sided minimal surfaces:
@@
See Theorem 3.6 below. We would like to point out that the uniqueness of the limit in geometric measure theory, in general, is quite hard when the multiplicity of convergence is greater than one.
Next, we further study what the possible limit minimal surfaces can be. Because the flow is the gradient of the area functional, one expects to typically find area minimizing or at least stable minimal surfaces using mean curvature flow. Nevertheless, in general, the long-time limit of mean curvature flow can be unstable. For curve shortening flow, a well-known example is the flow $\gamma_{t}$ of a closed embedded curve $\gamma\subset(S^{2},g_{\text{round}})$ which bisects the sphere. It will converge, as $t\to\infty$, smoothly to an equator of the sphere; in contrast, curves $\gamma$ which do not bisect the sphere flow to round points. Most curves $\gamma$ do not bisect the sphere, and we can always perturb one that bisects the sphere to one that does not. Therefore, most curve shortening flows cannot converge to an unstable geodesic. This observation motivates us to consider the extent to which we can show the flow will generically avoid unstable minimal surfaces.
The main tool we use to pursue this is the avoidance principle. In ${\mathbb{R}}^{n+1}$, if two mean curvature flows of closed hypersurfaces are disjoint at time 0, then they will remain disjoint for all future time. Moreover, the distance between these two flows is increasing. This basic property motivates the definition of weak set flows by Ilmanen [Ilmanen92_LSF] and White [White95_WSF_Top]. The avoidance principle in a manifold is more subtle, because in general two disjoint mean curvature flows can actually get closer – consider, for instance, the flow near an area minimizing minimal surface. In a manifold equipped with a positive Ricci curvature metric, the avoidance principle is very strong by White [White24_Avoidance], and lets us easily show the following:
@@
In a Ricci positive manifold $(N^{3},g)$, the mean curvature flow of a generic 22–sided embedded hypersurface $M$ will go extinct in finite time.
This can be viewed as a generalization of the generic behavior of the curve shortening flows in $S^{2}$. As a corollary, this yields another proof of the classical fact that $H_{2}(N,{\mathbb{R}})$ vanishes for 33–manifolds NN of positive Ricci curvature, see Corollary 4.1. In the case that $(N,g)$ is not Ricci positive there could be stable minimal surface, and we cannot expect the flows starting from generic embedded hypersurfaces to go extinct in finite time. However, we can use the avoidance principle to show that after perturbations, the flow can avoid those unstable limits. Our next theorem, inspired by the work of Colding and Minicozzi [ColdingMinicozzi12_generic], says that in closed 33–manifolds, from any embedded initial data, one may construct piecewise generic mean curvature flows which avoid unstable minimal surfaces as limits:
@@
Let $M^{2}$ be a 22–sided properly embedded surface of a closed compact Riemannian 33–manifold $(N^{3},g)$. Then there exists a piecewise almost regular flow emanating from it which either goes extinct in finite time, or there exist $t_{i}\to\infty$, such that $M_{t_{i}}$ converges possibly with multiplicity to a (potentially disconnected) stable minimal surface.
When the ambient metric is bumpy, which is a generic condition by work of White [White91_Bumpy, White17_Bumpy], we observe that the limit in the statement above is unique by Theorem 1.2. Note that in light of [ChenSun24_Multi2inManifold], there are examples where the flow asymptotically converges to a stable limiting surface with multiplicity greater than one, so higher multiplicity convergence in the above may actually occur, and in the case the limit is strictly stable seems to be robust under further perturbations. We will give the definition of piecewise almost regular flow in the section below, but roughly speaking, it is a concatenation of almost regular flows where, by hand, isotopies are performed at the jumps; the number of perturbations by isotopy taken will be finite, and the perturbations can be taken as small as we wish. In that sense, the statement above is a statement on generic flows following the convention of Colding and Minicozzi [ColdingMinicozzi12_generic].
Of course, it is certainly desirable to only require perturbing the initial data. The obstruction to doing so in the statement above essentially boils down to the possibility of convergence to an unstable minimal surface with multiplicity; in analogy to singularity analysis and the resolution of the Multiplicity One Conjecture by Bamler and Kleiner [BamlerKleiner23_Multiplicity1], one might expect this to not occur; see the analogous situation in Almgren-Pitts minmax theory [Zhou19_multi1]. We recall the following conjecture in [ChenSun24_Multi2inManifold]:
Conjecture 1.5 (Conjecture 1.3 in [ChenSun24_Multi2inManifold]).
@@
Despite this, we point out that as applications, we get flow proofs of the following existence theorems for stable minimal surfaces. In the proofs of these, we will often specifically use piecewise generic flows, because the topological change through singularities is easy to understand. We also recall that obtaining stability of the minimal surfaces is significant, because these often have better–controlled topology and geometry, for instance, in the context of positive scalar curvature. These corollaries are not new, and there have been long–standing proofs of them using geometric measure theory or harmonic map theory, see [FedererFleming60_Current, SchoenYau79_ExistenceIncompressibleMinSurf, SacksUhlenbeck81_minimal2sphere, MeeksYau80_Top3d, MeeksSimonYau82_MinSurf, FreedmanHassScott83_LeastArea], among others. On the other hand, the construction of these minimal surfaces via flows, along with a number of fairly explicit perturbations, is quite amenable to numerical approximation with the caveat that one has found appropriate initial data to start with. In the following statements $(N,g)$ is a compact orientable Riemannian 33–manifold, where in this case oriented surfaces are 22–sided and vice versa:
@@
Note that in the statement above, we invoke the sphere theorem to find embedded initial data for which to start the flow, and so the above does not constitute a new proof of it. As we pointed out already, compared to other techniques, a pitfall of using the mean curvature flow to find minimal surfaces is that it is best to start with well-prepared (i.e., embedded) initial data. In some cases, the flow can produce stable minimal surfaces of nontrivial topology, the most basic of which is the following.
@@
Suppose $N$ has a trivial second homotopy group and contains a homologically nontrivial embedded torus. Then $N$ contains an embedded stable minimal surface homeomorphic either to a torus or a Klein bottle.
Recall that hyperbolic manifolds are aspherical, so the space of such 33–manifolds with $\pi_{2}$ trivial is quite large and is even generic amongst 33–manifolds from a number of perspectives, for example, see [DunfieldTHurston06_Random3Mfd, Maher10_RandomHeegaard]. For surfaces of more general topology, we can show the following result, which is very much in the spirit of the existence result of Freedman, Hass, and Scott [FreedmanHassScott83_LeastArea] – in fact, it implies the result above as we discuss in Section 5. The surfaces in the Corollary below are known as algebraically incompressible surfaces, and are incompressible in the more geometric sense via Dehn’s lemma:
@@
By double cover of an embedded surface $P\subset N$, here we mean specifically a surface which is the boundary of a tubular neighborhood of $P$ in $N$. It is a fact (see, for instance, Lemma 3.6 in [Hatcher07_Notes3Mfd]) that every class of $H_{2}(N,{\mathbb{Z}})$ can be represented by an orientable properly embedded surface $M$. Recalling the Thurston norm $\sigma||_{T}$ of $\sigma\in H_{2}(N,{\mathbb{Z}})$ is the smallest attained genus of embedded surfaces homologous to $\sigma$. We see any such surface $M$ realizing the norm must be algebraically incompressible. If $\sigma||_{T}\geq 1$ and if we suppose that $\pi_{2}(N)$ is trivial, we can apply the Corollary above, and claim that the minimal surface we obtain realizes the Thurston norm. We can record the discussion above as follows:
@@
Suppose $\sigma\in H_{2}(N,{\mathbb{Z}})$, with $\pi_{2}(N)$ trivial. Then either there exists an embedded oriented stable minimal surface $\Sigma$ in the class of $\sigma$ that realizes the Thurston norm of $\sigma$ or there is an element realizing the Thurston norm which is the double cover of an embedded nonoriented stable minimal surface.
Finally, we note that Theorem 1.1 lets us substitute the tightening process in minmax for mean curvature flow, at least in some cases. A major issue in applying the flow to a sweepout is that the flow of a sweepout may not remain a sweepout, but this can be compensated for if the flows are not nonfattening. In particular, we show the following:
@@
Originally, this result was proved by Simon and Smith using minmax theory, see also [ColdingDeLellis03_minmax]. As to the actual best result known, we point out that recently Haslhofer and Ketover in [HaslhoferKetover19_Min2Sphere] showed there are at least two minimal spheres in $S^{3}$ with a bumpy metric, and very recently Zhichao Wang and Xin Zhou in [WangZhou23_4Sphere] proved that there are at least $4$ embedded minimal spheres in $S^{3}$ equipped with either a bumpy or Ricci positive metric. We expect that if Conjecture 1.5 is true, then the mean curvature flow can also be used to recover their results. We emphasize that in the above, we are not applying Theorem 1.4 and the minimal surface above is potentially unstable, as must be the case when, for instance, $N$ is Ricci positive.
We conclude the introduction with a remark on the compactness assumptions made throughout. When the ambient manifold is not closed, but merely complete with uniform bounded curvature and injectivity radius lower bound, Theorems 1.1 and 1.4 remain valid, if we include the possibility that the flow can escape from any compact set. For example, if $N$ is $S^{2}\times{\mathbb{R}}$ with an infinite long trumpet shape (i.e. a cusp) in one direction, the flow of an appropriate $S^{2}$ section of it will escape to spatial infinity/clear out, neither terminating nor converging to a minimal surface. It seems to be a more complicated manner yet to generalize our results to the case $M$ is allowed to be noncompact (say properly embedded in a noncompact $N$), one very basic reason being that the area can be infinite. As before, the flow may clear out in the manner of the example above, but it seems that in some cases, if the flow doesn’t clear out or go extinct, one could show it must flow to a minimal surface. For a basic example, a well known consequence of Ecker and Huisken [EckerHuisken89_EntireGraph] says that uniformly Lipschitz graphs of bounded height in ${\mathbb{R}}^{n}$ must converge to flat, hence minimal, planes.